If the numerator of the fraction is increased by 1 the fraction be comes 1if its numerator is increased by 9 it becomes 2.what is the fraction?

Grade:8

1 Answers

Arun
6 years ago
Dear Zia
 
let fraction is x/y
 
now,
(x+1)/y = 1
x+1 = y                    …...(1)
 
and also,
(x+9)/y = 2
x +9 = 2y                …...(2)
 
from (1) & (2)-
x+9 = 2(x+1)
x+9 = 2x + 2
x = 7
hence 
y = 8
 
hence fraction is 7/8
 
Regards
Arun (askIITians forum expert)